The standard deviation of lifetime of electric light bulbs is 100 hours. Find the 95% confidence limits for the population standard deviation for such electric bulbs.
1
Expert's answer
2020-10-14T18:23:47-0400

Confidence intervals for the true standard deviation can be constructed using the chi-square distribution.

if the original population of data is normally distributed, then the expression "\dfrac{s^2(n-1)}{\sigma^2}" has a chi-square distribution with  "n-1" degrees of freedom.

The 100(1−α)% confidence intervals

Two-sided "(1-\alpha)\%" confidence interval for "\sigma" is


"\dfrac{s\sqrt{n-1}}{\chi^2_{\alpha/2,n-1}}\leq \sigma\leq\dfrac{s\sqrt{n-1}}{\chi^2_{1-\alpha/2,n-1}}"

The sample must be taken from a normally distributed population.

If "\alpha=0.05, n=50"

"\chi^2_{L}=\chi^2_{1-\alpha/2,n-1}=32.3574"

"\chi^2_{R}=\chi^2_{\alpha/2,n-1}=71.4202"

Given "s=100"

Two-sided "95\%" confidence interval for "\sigma" is


"\dfrac{100\sqrt{50-1}}{71.4202}\leq \sigma\leq\dfrac{100\sqrt{50-1}}{32.3574}"



"9.8011\leq \sigma\leq21.6334"